OpenStudy (anonymous):

What are the sine, cosine, and tangent of 5 pi over 4 radians?

OpenStudy (anonymous):

please help

ganeshie8 (ganeshie8):

sin(5pi/4) = sin(pi + pi/4) = -sin(pi/4) = \(\large -\dfrac{\sqrt{2}}{2}\)

OpenStudy (anonymous):

tangent?

OpenStudy (anonymous):

cosine?

ganeshie8 (ganeshie8):

work cosine same way

ganeshie8 (ganeshie8):

cos(5pi/4) = cos(pi + pi/4) = -cos(pi/4) = \(\large -\dfrac{\sqrt{2}}{2}\)

OpenStudy (anonymous):

can you walk me through it please

ganeshie8 (ganeshie8):

so sin and cos have same values for 5pi/4 ^

ganeshie8 (ganeshie8):

For tan, simply take the ratio of sin and cos - you would get : tan(5pi/4) = \(\large 1\)

OpenStudy (anonymous):

sin Θ = negative square root 2 over 2; cos Θ = negative square root 2 over 2; tan Θ = 1

OpenStudy (anonymous):

right?

ganeshie8 (ganeshie8):

looks good

OpenStudy (anonymous):

thanks

ganeshie8 (ganeshie8):

yw